OpenStudy (anonymous):

what is the power dissapation for a series circuit that has 10V ,1.0 Kohms ,6.8 Kohms

OpenStudy (stormfire1):

In series current is constant and you can just add up the resistances to get 7.8 Kohms First, find the current in the circuit using the formula E = I * R: 10V / 7800ohms = .0013A Now find the total power dissipated using P = I * E P = .0013A * 7800 = 10.14W

OpenStudy (anonymous):

power dissipation is I^2 * R

OpenStudy (anonymous):

first take sqaure of 0.0013 and then multiply with 7800

OpenStudy (stormfire1):

Oops...I put the formula down but didn't apply it right. It should have been .0013A * 10V...which = .013W. The I^2*R formula does the same thing.